# Stellar Mass Segregation in Dark Matter Halos

### v1

@@ Abstract
We study the effect of stellar mass segregation driven by collisional relaxation within the potential well of a smooth dark matter halo. This effect is of particular relevance for old stellar systems with short crossing times, where small collisional perturbations accumulate over many dynamical time scales. We run collisional $N$-body simulations tailored to the ambiguous stellar systems Ursa Major 3/Unions 1, Delve 1 and Eridanus 3, modelling their stellar populations as two-component systems of high- and low-mass stars, respectively. For Ursa Major 3/Unions 1 (Delve 1), assuming a dynamical-to-stellar mass ratio of 10, we find that after 10 Gyr of evolution, the radial extent of its low-mass stars will be twice as large (40 per cent larger) than that of its high-mass stars. We show that weak tides do not alter this relative separation of half-light radii, whereas for the case of strong tidal fields, mass segregation facilitates the tidal stripping of low-mass stars. We further find that as the population of high-mass stars contracts and cools, the number of dynamically formed binaries within that population increases. Our results call for caution when using stellar mass segregation as a criterion to separate star clusters from dwarf galaxies, and suggest that mass segregation increases the abundance of massive binaries in the central regions of dark matter-dominated dwarf galaxies.
@@ 1 Introduction
In cold dark matter cosmology, galaxies are expected to form deep within the potential wells of dark matter halos (White & Rees, 1978). Numerical simulations suggest that these cold dark matter halos reach remarkably high central densities, well-described by a universal centrally-divergent density profile (Navarro et al., 1996, 1997). The centrally-divergent centers of cold dark matter halos are commonly referred to as “cusps”. Density cusps render cold dark matter halos resilient to the effect of tides (Peñarrubia et al., 2010; van den Bosch et al., 2018): for a fixed tidal field strength, it is argued that cold dark matter halos cannot be tidally stripped beyond a certain point and instead converge towards a stable asymptotic remnant state (Errani & Navarro, 2021; Stücker et al., 2023). Stars embedded in such cold dark matter cusps would be protected from tidal disruption, plausibly giving rise to a population of “micro galaxies” (Errani & Peñarrubia, 2020; Errani et al., 2024a). The discovery of such objects would allow to put strong constraints on the nature of dark matter, as discussed in the context of a potential self-annihilation signal (Crnogorčević & Linden, 2024; Errani et al., 2024b), primordial black hole dark matter (Graham & Ramani, 2024) or ultra-light particle dark matter (Safarzadeh & Spergel, 2020).
@@
In recent years, deep photometric surveys have led to the discovery of several objects with structural parameters at the interface of the globular cluster- and dwarf galaxy regimes (Conn et al., 2018; Mau et al., 2020; Cerny et al., 2023a, b; Smith et al., 2024). Could some of these systems possibly be among the faintest yet dark matter-dominated dwarf galaxies known to date (Errani et al., 2024b; Smith et al., 2024; Simon et al., 2024)? Proving unequivocally that these objects are either dark matter-dominated dwarf galaxies or star clusters devoid of any dark matter has been proven a highly challenging task.
@@
Traditionally, stellar kinematics have been used to argue in favor of the high dark matter content of dwarf galaxies (Mateo, 1998; Walker et al., 2007). For faint stellar systems with few member stars, the inferred velocity dispersions are shown to sensitively depend on the inclusion- or exclusion of individual members stars (Smith et al., 2024) and the choice of prior (Simon et al., 2024). The presence of binary stars further complicates such studies by adding a velocity dispersion floor that, in particular for low-mass dwarf galaxies, needs to be accurately accounted for and generally requires the availability of multi-epoch spectroscopic measurements (McConnachie & Côté, 2010; Buttry et al., 2022).
@@
Another pathway suggested to distinguish dwarf galaxies from globular clusters has been to use the element abundance patterns of their stars (see e.g. Gratton et al. 2012, Bastian & Lardo 2018 for a discussion of element abundances in globular clusters, and Venn et al. 2004, Ji et al. 2019 for dwarf galaxies), though their application to faint stellar systems with few member stars remains challenging (Zaremba et al., 2025).
@@
A further strategy discussed in the literature is based on tidal survival: the existence of the ancient stellar system Ursa Major 3/Unions 1 (Smith et al., 2024) in the inner region of the Milky Way has been used to argue in favour of it being embedded in a dark matter halo and in turn protected from tidal disruption (Errani et al., 2024b). This picture though has been recently challenged by Devlin et al. (2025) who argue that the baryonic mass of Ursa Major 3/Unions 1 has been underestimated in its discovery paper, making it more stable against Milky Way tides even in absence of a surrounding dark matter halo.
@@
The paper is structured as follows. In section 2, we estimate the time scale for collisional relaxation in presence of a smooth dark matter halo. In section 3, we detail the numerical setup of our collisional $N$-body simulations. In section 4.1, we discuss the results of our simulations for the case of a static dark matter halo surrounding each stellar system, and describe the dynamical formation of stellar binaries in Sec. 4.2. We extend our analysis to include the effects of tides through a time-evolving dark matter potential in section 4.4. Finally, we summarize our results and conclusions in section 5.
@@ 2 Relaxation Times
The stellar component has a (3D) velocity dispersion of roughly $\langle v^{2}\rangle\approx GM_{\mathrm{h}}/r_{\mathrm{h}}=GM_{\star}\Upsilon_{\mathrm{dyn}}/(2r_{\mathrm{h}})$, and a crossing time of
@@
For the case of UMa3/U1, for example, we find $T_{\mathrm{cross}}\approx 13\,\mathrm{Myr}$ when assuming a dynamical-to-stellar mass ratio $\Upsilon_{\mathrm{dyn}}=10$: over $10\,\mathrm{Gyr}$, the system has gone through close to ${\sim}800$ crossing times. Over one crossing time, the average squared velocity increase that an individual star experiences due to the fluctuating gravitational potential of $N_{\star}$ stars with individual masses $m_{\star}$ can be approximated by (using Eq. 18 in Peñarrubia 2019 and assuming that the stellar number density $\bar{n}=3N_{\star}/(8\pi r_{\mathrm{h}}^{3})$ is approximately constant within the half-light radius):
@@
$T_{\mathrm{rel}}$$\approx$ $1/\sqrt{12\pi G}\,\left(\Upsilon_{\mathrm{dyn}}r_{\mathrm{h}}\right)^{3/2}N_{\star}M_{\star}^{-1/2}\left[\ln(\Lambda)-1.9\right]^{-1}$ (5)
@@
All else being equal, the more dark-matter dominated the system is, the larger is its relaxation time.
@@
Figure 1: Several Milky Way satellites have relaxation times $T_{\mathrm{rel}}$ smaller than $10\,\mathrm{Gyr}$ even in presence of substantial amounts of dark matter. Those objects may exhibit stellar mass segregation driven by collisional relaxation. Shown are stellar masses $M_{\star}$ and (3D) half-light radii $r_{\mathrm{h}}$ of dwarf galaxies (open circles), ambiguous stellar clusters (triangles), as well as the “micro galaxy” candidates UMa3/U1, Del1 and Eri3 (red filled circle, diamond and square, respectively). See footnote 4 for references. Curves of constant relaxation time $T_{\mathrm{rel}}=1\,\mathrm{Gyr}$ and $10\,\mathrm{Gyr}$ are computed assuming a dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$ within the half-light radius, with stellar masses sampled from a Chabrier (2003) present-day mass function (see text for details).
In Fig. 1, we show Relaxation times computed by summing Eq. 4 over individual stellar masses $m_{\star}$ drawn from a Chabrier (2003) present-day mass function (PDMF), together with a compilation [^4] Properties of faint stellar clusters are as compiled in Cerny et al. (2023a, b). The dwarf galaxy data is taken from McConnachie (2012), version January 2021, with updates for Bootes 2 (Bruce et al., 2023), Draco 2 (Martin et al., 2016a; Longeard et al., 2018) and TriII (Martin et al., 2016b; Kirby et al., 2017). For the faint stellar systems UMa3/U1 (Smith et al., 2024), Del1 (Mau et al., 2020) and Eri3 (Conn et al., 2018) see Table 1 and footnotes 1, 1, 1. of stellar masses $M_{\star}$ and (3D) half-light radii $r_{\mathrm{h}}$ of Milky Way dwarf galaxies and faint clusters with structural properties at the interface of the globular cluster- and dwarf galaxy regimes. In Fig. 1, we assume $\Upsilon_{\mathrm{dyn}}=10$, but the results are easily translated to other choices for dynamical-to-stellar mass ratio through Eq. 5. Uncertainties on the relaxation times shown here stem from sampling the stellar mass function and span the $16^{\mathrm{th}}$ to $84^{\mathrm{th}}$ percentile of crossing times for random realizations of total mass $M_{\star}$.
For the faint stellar systems UMa3/U1 and Delve1, assuming a dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$, we find relaxation times of $0.9$ and $8.4\,\mathrm{Gyr}$, respectively, substantially lower than the age of their stellar populations. For these systems, we can therefore expect dynamical signatures of collisional processes, such as stellar mass segregation, even when embedded in a smooth, static and gravitationally dominant dark matter subhalo.
@@ 3.1 Example systems
We build our $N$-body models to approximately resemble the faint stellar systems UMa3/U1 (Smith et al., 2024), Del1 (Mau et al., 2020) and Eri3 (Conn et al., 2018). These systems are plausible candidates for being the smallest dwarf galaxies known to date, motivated by their kinematics and chemistry (Simon et al., 2024), or by their tidal survival on a low-pericentre orbit (Errani et al., 2024b). Table 1 lists the structural parameters used for our $N$-body models. We estimate the (3D) half-light radii $r_{\mathrm{h}}\approx 4R_{\mathrm{h}}/3$ from the the published projected radii $R_{\mathrm{h}}$. For UMa3/U1 and Del1, we use total stellar masses as published; for Eri3, we estimate the stellar stellar mass from its luminosity, assuming a stellar mass-to-light ratio of $\Upsilon_{\star}=1.4$. This value matches the stellar mass-to-light ratio inferred in Smith et al. (2024) and Mau et al. (2020) for UMa3/U1 and Del1, respectively.
Table 1: Parameters used for the $N$-body models. The table lists the total stellar mass $M_{\star}$ (for the case of Eri3 estimated from its luminosity), the (3D) half-light radius $r_{\mathrm{h}}\approx 4R_{\mathrm{h}}/3$ estimated from the projected $R_{\mathrm{h}}$, and the total number of star particles $N_{\star}$. For reference, we also list the resulting crossing- (Eq. 2) and relaxation times (Eq. 5) assuming $\Upsilon_{\mathrm{dyn}}=10$ and $20$.

### v2

@@ Abstract
We study the effect of stellar mass segregation driven by collisional relaxation within the potential well of a smooth dark matter halo. This effect is of particular relevance for old stellar systems with short crossing times, where small collisional perturbations accumulate over many dynamical timescales. We run collisional $N$-body simulations tailored to the ambiguous stellar systems Ursa Major 3/Unions 1, Delve 1 and Eridanus 3, modeling their stellar populations as two-component systems of high- and low-mass stars, respectively. For Ursa Major 3/Unions 1 (Delve 1), assuming a dynamical-to-stellar mass ratio of 10, we find that after 10 Gyr of evolution, the radial extent of its low-mass stars will be twice as large as (40 per cent larger than) that of its high-mass stars. We show that weak tides do not alter this relative separation of half-light radii, whereas for the case of strong tidal fields, mass segregation facilitates the tidal stripping of low-mass stars. We further find that as the population of high-mass stars contracts and cools, the number of dynamically formed binaries within that population increases. Our results call for caution when using stellar mass segregation as a criterion to separate star clusters from dwarf galaxies, and suggest that mass segregation increases the abundance of massive binaries in the central regions of dark matter-dominated dwarf galaxies.
@@ 1 Introduction
In cold dark matter cosmology, galaxies are expected to form deep within the potential wells of dark matter halos (White & Rees, 1978). Numerical simulations suggest that these cold dark matter halos reach remarkably high central densities, well-described by a universal centrally-divergent density profile (Navarro et al., 1996, 1997). The centrally-divergent centers of cold dark matter halos are commonly referred to as “cusps”. Density cusps render cold dark matter halos resilient to the effect of tides (Peñarrubia et al., 2010; van den Bosch et al., 2018): for a fixed tidal field strength, it is argued that cold dark matter halos cannot be tidally stripped beyond a certain point and instead converge toward a stable asymptotic remnant state (Errani & Navarro, 2021; Stücker et al., 2023). Stars embedded in such cold dark matter cusps would be protected from tidal disruption, plausibly giving rise to a population of “micro galaxies” (Errani & Peñarrubia, 2020; Errani et al., 2024a). The discovery of such objects would allow us to put strong constraints on the nature of dark matter, as discussed in the context of a potential self-annihilation signal (Crnogorčević & Linden, 2024; Errani et al., 2024b), primordial black hole dark matter (Graham & Ramani, 2024) or ultra-light particle dark matter (Safarzadeh & Spergel, 2020).
@@
In recent years, deep photometric surveys have led to the discovery of several objects with structural parameters at the interface of the globular cluster and dwarf galaxy regimes (Conn et al., 2018; Mau et al., 2020; Cerny et al., 2023a, b; Smith et al., 2024). Could some of these systems possibly be among the faintest dark matter-dominated dwarf galaxies known to date (Errani et al., 2024b; Smith et al., 2024; Simon et al., 2024)? Proving unequivocally that these objects are either dark matter-dominated dwarf galaxies or star clusters devoid of any dark matter has turned out to be a very challenging task.
@@
Traditionally, stellar kinematics have been used to argue in favor of the high dark matter content of dwarf galaxies (Mateo, 1998; Walker et al., 2007). For faint stellar systems with few member stars, the inferred velocity dispersions are shown to depend sensitively on the inclusion or exclusion of individual members stars (Smith et al., 2024) and the choice of prior (Simon et al., 2024). The presence of binary stars further complicates such studies by adding a velocity dispersion floor that, particularly for low-mass dwarf galaxies, needs to be accurately accounted for and generally requires the availability of multi-epoch spectroscopic measurements (McConnachie & Côté, 2010; Buttry et al., 2022).
@@
Another pathway suggested to distinguish dwarf galaxies from globular clusters has been to use the elemental abundance patterns of their stars (see, e.g., Gratton et al. 2012, Bastian & Lardo 2018 for a discussion of elemental abundances in globular clusters, and Venn et al. 2004, Ji et al. 2019 for dwarf galaxies), though their application to faint stellar systems with few member stars remains challenging (see, e.g., Zaremba et al. 2025; see also Fu et al. 2023 who use narrow-band imaging to infer metallicity dispersions for faint stellar systems where spectroscopic measurements are not feasible).
@@
A further strategy discussed in the literature is based on tidal survival: the existence of the ancient stellar system Ursa Major 3/Unions 1 (Smith et al., 2024) in the inner region of the Milky Way has been used to argue in favour of it being embedded in a dark matter halo and in turn protected from tidal disruption (Errani et al., 2024b). This picture, though, has recently been challenged by Devlin et al. (2025), who argue that the baryonic mass of Ursa Major 3/Unions 1 has been underestimated in its discovery paper, making it more stable against Milky Way tides even in the absence of a surrounding dark matter halo.
@@
The paper is structured as follows. In section 2, we estimate the timescale for collisional relaxation in the presence of a smooth and static dark matter halo. In section 3, we detail the numerical setup of our collisional $N$-body simulations. In section 4.1, we discuss the results of our simulations for the case of a static dark matter halo surrounding each stellar system, and in Sec. 4.2 we describe the dynamical formation of stellar binaries. We extend our analysis to a broader range of initial dynamical-to-stellar mass ratios in Sec. 4.3. In Sec. 4.4, we study the effect of galactic tides on stellar mass segregation through a time-evolving dark matter potential. Finally, we summarize our results and conclusions in section 5.
@@ 2 Relaxation Times
The dynamical-to-stellar mass ratio $\Upsilon_{\mathrm{dyn}}$ can be observationally constrained via the line-of-sight velocity dispersion $\sigma_{\mathrm{los}}$ using the relation $\Upsilon_{\mathrm{dyn}}\approx 8R_{\mathrm{h}}\sigma^{2}_{\mathrm{los}}G^{-1}M_{\star}^{-1}$, where $R_{\mathrm{h}}$ is the projected half-light radius (see Eq. 8 of Errani et al. 2024b). For the systems UMa3/U1, Del1, and Eri3, the upper limits on their respective velocity dispersions (as listed in Sec. 1) yield approximate constraints of $\Upsilon_{\mathrm{dyn}}\lesssim 5600$, $120$ and $580$, respectively. If these objects are indeed dark matter-dominated relics of the Milky Way’s early accretion history, their true dynamical-to-stellar mass ratios could be significantly lower than the current upper bounds. Taking UMa3/U1 as an example, its pericentre and apocentre of ${\sim}13\,\mathrm{kpc}$ and ${\sim}26\,\mathrm{kpc}$ (Smith et al., 2024), respectively, suggest an early accretion onto the Milky Way (see Fig. 1 of Errani et al. 2024a). This makes it plausible that the subhalo surrounding UMa3/U1 is structurally close to the maximally stripped “tidal remnant” state for its orbit (Errani & Navarro, 2021), corresponding to a mean enclosed dark matter density of order $\bar{\rho}_{\mathrm{h}}\sim 10^{9}\,\mathrm{M_{\odot}}\,\mathrm{kpc}^{-3}$ and a dynamical-to-stellar mass ratio of order $\Upsilon_{\mathrm{dyn}}\sim 30$ (see Fig. 4 and 5 of Errani et al. 2024b). As a working example, we adopt a fiducial [^2] Note that the total mass-to-luminosity ratio $M_{\mathrm{tot}}/L$, expressed in units of $\mathrm{M_{\odot}}/\mathrm{L}_{\odot}$, may be substantially higher than the value of $\Upsilon_{\mathrm{dyn}}$, and depends sensitively on how much the dark matter subhalo extends beyond the stellar half-light radius. value of $\Upsilon_{\mathrm{dyn}}=10$ and later explore a broader range, $3\lesssim\Upsilon_{\mathrm{dyn}}\lesssim 300$, in Sec. 4.3.
A stellar population that is embedded in a dark matter subhalo has a (3D) velocity dispersion of roughly $\langle v^{2}\rangle\approx GM_{\mathrm{h}}/r_{\mathrm{h}}=GM_{\star}\Upsilon_{\mathrm{dyn}}/(2r_{\mathrm{h}})$, and a crossing time of
@@
For the case of UMa3/U1, for example, we find $T_{\mathrm{cross}}\approx 13\,\mathrm{Myr}$ when assuming a dynamical-to-stellar mass ratio $\Upsilon_{\mathrm{dyn}}=10$: over $10\,\mathrm{Gyr}$, the system goes through close to ${\sim}800$ crossing times. Over one crossing time, the average increase in squared velocity that an individual star experiences due to the fluctuating gravitational potential of $N_{\star}$ stars with individual masses $m_{\star}$ can be approximated by (using Eq. 18 of Peñarrubia 2019 and assuming that the stellar number density $\bar{n}=3N_{\star}/(8\pi r_{\mathrm{h}}^{3})$ is approximately constant within the half-light radius)
@@
$T_{\mathrm{rel}}$ $\approx$ $\!\bigl(1/\sqrt{12\pi G}\bigr)\left(\Upsilon_{\mathrm{dyn}}r_{\mathrm{h}}\right)^{3/2}N_{\star}M_{\star}^{-1/2}\left[\ln(\Lambda)-1.9\right]^{-1}$ (5)
@@
All else being equal, the more dark matter-dominated the system is, the longer is its relaxation time.
@@
Figure 1: Several Milky Way satellites have relaxation times $T_{\mathrm{rel}}$ shorter than $10\,\mathrm{Gyr}$ even in the presence of substantial amounts of dark matter. Those objects may exhibit stellar mass segregation driven by collisional relaxation. Shown are stellar masses $M_{\star}$ and (3D) half-light radii $r_{\mathrm{h}}$ of dwarf galaxies (open circles), ambiguous stellar clusters (triangles), as well as the “micro galaxy” candidates UMa3/U1, Del1 and Eri3 (red filled circle, diamond and square, respectively). See footnote 5 for references. Curves of constant relaxation time $T_{\mathrm{rel}}=1\,\mathrm{Gyr}$ and $10\,\mathrm{Gyr}$ are computed assuming a dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$ within the half-light radius, with stellar masses sampled from a Chabrier (2003) present-day mass function (see text for details).
In Fig. 1, we show relaxation times computed by summing Eq. 4 over individual stellar masses $m_{\star}$ drawn from a Chabrier (2003) present-day mass function (PDMF), together with a compilation [^5] Properties of faint stellar clusters are as compiled in Cerny et al. (2023a, b). The dwarf galaxy data is taken from McConnachie (2012), version January 2021, with updates for Boötes 2 (Bruce et al., 2023), Draco 2 (Martin et al., 2016a; Longeard et al., 2018) and TriII (Martin et al., 2016b; Kirby et al., 2017). For the faint stellar systems UMa3/U1 (Smith et al., 2024), Del1 (Mau et al., 2020) and Eri3 (Conn et al., 2018) see Table 1 and footnotes 1, 1, 1. of stellar masses $M_{\star}$ and (3D) half-light radii $r_{\mathrm{h}}$ of Milky Way dwarf galaxies and faint clusters with structural properties at the interface of the globular cluster and dwarf galaxy regimes. In Fig. 1, we assume $\Upsilon_{\mathrm{dyn}}=10$, but the results are easily translated to other choices for dynamical-to-stellar mass ratio through Eq. 5. Uncertainties on the relaxation times shown here stem from sampling the stellar mass function and span the $16^{\mathrm{th}}$ to $84^{\mathrm{th}}$ percentiles of crossing times for random realizations of total mass $M_{\star}$.
For the faint stellar systems UMa3/U1 and Delve1, assuming a dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$, we find relaxation times of $0.9$ and $8.4\,\mathrm{Gyr}$, respectively, substantially shorter than the age of their stellar populations. For these systems, we can therefore expect dynamical signatures of collisional processes, such as stellar mass segregation, even when embedded in a smooth, static and gravitationally dominant dark matter subhalo.
@@ 3.1 Example systems
We build our $N$-body models to approximately resemble the faint stellar systems UMa3/U1 (Smith et al., 2024), Del1 (Mau et al., 2020) and Eri3 (Conn et al., 2018). These systems are plausible candidates for being the smallest dwarf galaxies known to date, motivated by their kinematics and chemistry (Simon et al., 2024), or by their tidal survival on a low-pericentre orbit (Errani et al., 2024b). Table 1 lists the structural parameters used for our $N$-body models. We estimate the (3D) half-light radii $r_{\mathrm{h}}\approx 4R_{\mathrm{h}}/3$ from the published projected radii $R_{\mathrm{h}}$. For UMa3/U1 and Del1, we use total stellar masses as published; for Eri3, we estimate the stellar mass from its luminosity, assuming a stellar mass-to-light ratio of $\Upsilon_{\star}=1.4$. This value matches the stellar mass-to-light ratio inferred by Smith et al. (2024) and Mau et al. (2020) for UMa3/U1 and Del1, respectively.
Table 1: Parameters used for the $N$-body models. The table lists the total stellar mass $M_{\star}$ (for the case of Eri3 estimated from its luminosity), the (3D) half-light radius $r_{\mathrm{h}}\approx 4R_{\mathrm{h}}/3$ estimated from the projected $R_{\mathrm{h}}$, and the total number of star particles $N_{\star}$. For reference, we also list the resulting crossing (Eq. 2) and relaxation times (Eq. 5) assuming $\Upsilon_{\mathrm{dyn}}=10$ and 2020.
@@
Figure 2: Stellar mass segregation in NN-body realizations of the example systems UMa3/U1 (top), Del1 (center) and Eri1 (bottom). Each system is embedded in a cuspy dark matter halo, with an initial dynamical-to-stellar mass ratio of $\Upsilon_{\mathrm{dyn}}=10$. Shown are $\{x,y\}$ projections of snapshots at $t/\mathrm{Gyr}=0$, $2.5$, 55, $7.5$ and $10$. The axes are expressed in units of the initial half-light radius $r_{\mathrm{h0}}$. Individual high-mass stars are shown as dark grey points, while low-mass stars are shown in blue. The median half-light radii of high-mass and low-mass stars (computed over a sample of $N$-body realizations, see text) are shown as black and blue circles, respectively. Over a period of $10\,\mathrm{Gyr}$, the half-light radius of the population of low-mass stars expands, while the half-light radius of the high-mass population contracts. Consistent with the relaxation time estimates of Fig. 1, the effect is largest for the UMa3/U1 model, and smallest for the Eri3 model. Dynamically formed binaries consisting of two high-mass stars are highlighted in red (see Sec. 4.2). A video supplementing this figure is available in Appendix A.